Green Test Laboratory (GTL) performs heat testing and stress testing on materials and operates at full capacity. Under its current simple costing system, GTL aggregates all operating costs of $2,404,000 into a single cost pool. GTL calculates a rate per test-hour of $18.78, ($2,404,000/128,000 total tests-hours). Heat testing uses 80,000 test-hours and stress testing uses 48,000 test-hours. Lee Chang, GTL’s management accountant, believes that there is enough variation in test procedures and cost structures to establish separate costing and billing rates for heat testing and stress testing. The market for test services is becoming increasingly competitive. Without this information, any incorrect costing and pricing of its services could cause GTL to lose business. Lee divides GTL’s costs into four activity cost categories as follows:

  1. Direct labour costs, $450,000. These costs can be directly traced to heat testing, $350,000, and to stress testing, $100,000.
  2. Equipment related costs (e.g. rent, maintenance, energy), $750,000. These costs are allocated to heat and stress testing on the basis of test hours.
  3. Set up costs, $700,000. These costs are allocated to heat and stress testing on the basis of the number of set up hours required. Heat testing requires 13,500 set up hours and stress testing requires 4,000 set up hours.
  4. Costs of designing tests, $504,000. These costs are allocated to heat and stress testing on the basis of the time required to design the tests. Heat testing requires 4,200 design hours, and stress testing requires 2,100 design hours.

REQUIRED:

  1. Classify each activity cost as either: output unit level cost, batch level cost, product or service sustaining cost, or organisation sustaining cost. Justify the classification of each activity cost.

  1. Calculate the cost per test hour for heat testing and for stress testing.

  1. Explain how GTL’s management could use the cost hierarchy and Activity Based Costing (ABC) information to manage its business better.
  1. ’If you increase the number of indirect-cost pools in the company’s costing system, you will definitely increase the accuracy of the costs of outputs.’ Do you agree or disagree? Explain your answer in two points.
Answer #1

Question A

1. Direct Related Costs - Output Unit Level Cost

2. Equipment Related Costs - Output Unit Level Cost

3. Set-up Costs - Batch Level Cost

4. Cost of Designing Tests - Service Sustaining Costs

Notes

For 1 and 2 - Both Direct Labour Cost and Equipment Related Costs are both directly traceable according to the number of tests conducted that's why they are classified as Output Unit Level Cost.

For 3 - Setup Costs are the cost which are usually incurred after a certain lot or batch is performed they are least affected by each unit so they are Batch Level Cost.

For 4 - This is a cost which is neither batch related nor unit related. It is a cost for the organisation as a whole.

Requirement 2

Particulars Heat Testing Stress Testing
Direct Labour Cost 350,000 100,000
Equipment Related Cost 468,750 281,250
Setup Costs 540,000 160,000
Cost of Designing Tests 336,000 168,000
Total Costs 1,694,750 709,250
÷ Total Test Hour 80,000 48,000
Per Test Cost 21.18 14.78

NOTES

Direct Labour Cost are directly given in the question itself.

Equipment Related Cost = 750,000 / 128,000 Test Hour = $ 5.86 per Test Hour

For Heat Testing = 80,000 Test Hours * $ 5.86 Test Hour = $ 468,750

For Stress Testing = 48,000 Test Hour * $ 5.76 per Test Hour = $ 281,250

Setup Costs = 700,000 / 17,500 Set-up = $ 40 per Setup

For Heat Testing = 13,500 Setup * $ 40 per Setup = $ 540,000

For Stress Testing = 4,000 Setup * $ 40 per Setup = $ 160,000

Cost of Design Test = 504,000 / 6,300 = $ 80 per Design Hour

For Heat Testing = 4,200 Design Hour * $ 80 Per Hour = $ 336,000

For Stress Testing = 2,100 Design Hour * $ 80 per Design Hour = $ 168,000

Requirement 3

GTL can use the Cost Hierarchy and Activity Based Costing to manage its business better.

The reason behind the same is that with the traditional costing system the company was allocating the costs based on test hours which is not correct becuase the overall Overhead costs is not only affected by test hours as both heat test and stress test consume various type of activities and which also means consumption of various resources so how it is advisable cto allocate costs which is not relatable to the costs but with the help of Activity Based Costing there is a reasonable and more accurate distribution of Overhead costs which will help the GTL company in better pricing of test in the present competitive market and take an advantage position in the market.

Requirement 4

The Statement is totally correct. Following point will help in understanding

1. If the Indirect Costs pools are increased it will become more easy to track the various costs which are ignored earlier which will help to allocate the overhead more appropriately and the service/ product should get priced more accordingly.

2. More indirect cost pools means that it is easy to relate each cost with the consumption of resources used by the particular unit so it becomes easy to find the per unit cost which makes it easier to price that unit and helps in getting knowledge of the per unit income.